# Testing Stochastic Block Models Based on Maximum Sampling Entry-Wise Deviations

### v1

@@ 4 Real data analyses
To demonstrate the usefulness of our proposed method for both dense and sparse networks, we analyze two real datasets. Based on simulation studies, we find that our proposed method and Hu et al. (2021)’s method generally outperform those methods proposed by Lei (2016), Han et al. (2023), and Zhang and Amini (2023). Thus, in our real data analyses, we only consider our proposed method and Hu et al. (2021)’s method. The first dataset is the international trade data, which has been studied by Westveld and Hoff (2011), Lei (2016), and Hu et al. (2021). This data set contains yearly trade information of 58 countries. Hu et al. (2021) constructed a network with density of 48.1% (=1590/(58×57)×1590/(58$\times$ 57$\times$1590 / ( 58 × 57 ) × 100%), and it is a typical dense network. As noticed by Hu et al. (2021), the augmentation procedure may not work well for this data set because the size of the network is small and the added community may have less than or equal to two nodes. Therefore, we apply the test statistics, $T_{n,boot}$ and $T_{Hu,boot}$, to analyze the international trade network. The second data set is the co-authorship network dataset collected by Ji and Jin (2016). This network has 3,607 nodes and the density of this network is approximately equal to 0.012% (=1561/(3607×3606)×1561/(3607$\times$ 3606$\times$1561 / ( 3607 × 3606 ) ×100%), which makes it a sparse network. The simulation results indicate that $T^{+}_{n,boot}$ and $T^{+}_{Hu,boot}$ perform better than $T_{n}$ and $T_{Hu}$, respectively. Hence, we employ them to study the co-authorship network. However, the augmented tests are not applicable for testing $H_{0}:K_{0}=1$ since it is infeasible to specify the connection probability between communities. Thus, $T_{n,boot}$ and $T_{Hu,boot}$ are considered in this case. Finally, the number of bootstraps used in both examples is $M=100$, and the number of samples drawn from $n$ nodes is $B=\tilde{Q}^{-1/2}\{n/\log n\}^{1/3}$ as given in simulation studies.
@@ 4.1 Dense network case: International trade data
This dataset contains yearly trade information for $n=58$ countries from 1981 to 2000. For any given year, $\text{Trade}_{ij}$ represents the value of exports from country $i$ to country $j$; further descriptions can be found in Westveld and Hoff (2011). Following Saldana et al. (2017) and Hu et al. (2021), we consider the international trade data in 1995, and construct the network from the 58 countries based on $\text{Trade}_{ij}$ as follows. Each node represents a country, and define $W_{ij}=\text{Trade}_{ij}+\text{Trade}_{ji}$. Then, the adjacency matrix $A=(a_{ij})$ is defined such that $a_{ij}=1$ if $W_{ij}>W_{0.5}$ and $a_{ij}=0$ otherwise, where $W_{0.5}$ is the 50% quantile of the $W_{ij}$ values. As discussed in Saldana et al. (2017), it is possible to decompose the network into 3, 7, and 10 communities by three different methods. However, for different choice of $K$, the resulting communities from these 58 countries vary. Therefore, determining the appropriate number of communities is critical to label communities and make meaningful interpretations for 58 countries. To this end, we employ the statistics $T_{n,boot}$ and $T_{Hu,boot}$ to test the three hypotheses $H_{0}:K=3$, $H_{0}:K=7$, and $H_{0}:K=10$, respectively.
Given the 5% significance level, we apply our Theorem 1 and Hu et al. (2021)’s Theorem 1 to obtain the critical values of $T_{n,boot}$ and $T_{Hu,boot}$, which are $t_{n,0.95}=4.792$ and $t_{Hu,0.95}=3.413$, respectively. For $K_{0}=3,7,$ and 10, Hu et al. (2021) calculated their corresponding test statistics, $T_{Hu,boot}=44.28,3.33$ and 13.44. Comparing them with the critical value, 3.413, they concluded that $K=7$ is a reasonable choice. We next compute our proposed test statistics, $T_{n,boot}=15.02,3.07,$ and 16.76, for $K_{0}=3,7,$ and 10, respectively. Comparing them with the critical value, 4.792, we find that $T_{n,boot}$ does not reject the hypothesis $H_{0}:K=7$, whereas it rejects the hypotheses $H_{0}:K=3$ and $H_{0}:K=10$. Hence, $T_{n,boot}$ yields the same finding as that of Hu et al. (2021).
@@ 4.2 Sparse network case: Co-authorship data
Using this sparse network data, we test the hypothetical number of communities $K_{0}\in[1,8]$. Given the 5% significance level, we have $t_{n,0.95}=4.792$ and $t_{Hu,0.95}=3.413$ as given in the above example. In testing $H_{0}:K=1$, both $T_{n,boot}$ and $T_{Hu,boot}$ are larger than 400, which results in rejecting the null hypothesis. For $K_{0}\in[2,8]$, we have $T^{+}_{Hu,boot}>7$, which results in rejecting all of their corresponding hypotheses. This finding is not surprising since $T^{+}_{Hu,boot}$ fails to converge to an extreme-value distribution for sparse networks. In testing $K_{0}\in[2,8]$, we calculate $T^{+}_{n,boot}$=14.81, 15.80, 4.23, -1.53, 3.75, -2.52, and 3.08. As a result, $T^{+}_{n,boot}$ does not reject each of those hypotheses from $H_{0}:K=4$ to $H_{0}:K=8$ at the 5% significance level. Hence, $T^{+}_{n,boot}$ can identify $K=4$ to $K=8$ communities for all authors.
@@
In sum, our proposed test and Hu et al. (2021)’s test lead to the same result in the dense network example, and we also provide more interpretations based on the Editor’s comments. In the sparse network example, our test successfully detects a reasonable number of network communities, whereas Hu et al. (2021)’s test rejects all possible hypotheses and fails to yield any conclusion. Consequently, our proposed method is applicable for both data sets.
@@ 5 Concluding remarks
12013Agarwal et al.Agarwal et al. (2013)agarwal2013traffic Agarwal, S., Kodialam, M., and Lakshman, T. (2013), “Traffic Engineering in Software Defined Networks,” 2013 Proceedings IEEE INFOCOM, 2211–2219. 22013Amini et al.Amini et al. (2013)amini2013pseudo Amini, A. A., Chen, A., Bickel, P. J., and Levina, E. (2013), “Pseudo-Likelihood Methods for Community Detection in Large Sparse Networks,” The Annals of Statistics, 41, 2097–2122. 32009Bickel and ChenBickel and Chen (2009)bickel2009nonparametric Bickel, P. J., and Chen, A. (2009), “A Nonparametric View of Network Models and Newman–Girvan and Other Modularities,” Proceedings of the National Academy of Sciences, 106, 21068–21073. 42016Bickel and SarkarBickel and Sarkar (2016)bickel2016hypothesis Bickel, P. J., and Sarkar, P. (2016), “Hypothesis Testing for Automated Community Detection in Networks,” Journal of the Royal Statistical Society Series B: Statistical Methodology, 78, 253–273. 52015ChatterjeeChatterjee (2015)chatterjee2015matrix Chatterjee, S. (2015), “Matrix Estimation by Universal Singular Value Thresholding,” The Annals of Statistics, 43, 177–214. 62018Chen and LeiChen and Lei (2018)chen2018network Chen, K., and Lei, J. (2018), “Network Cross-Validation for Determining the Number of Communities in Network Data,” Journal of the American Statistical Association, 113, 241–251. 72015Chin et al.Chin et al. (2015)chin2015stochastic Chin, P., Rao, A., and Vu, V. (2015), “Stochastic Block Model and Community Detection in Sparse Graphs: A Spectral Algorithm with Optimal Rate of Recovery,” Conference on Learning Theory, 391–423. 82012Choi et al.Choi et al. (2012)choi2012stochastic Choi, D. S., Wolfe, P. J., and Airoldi, E. M. (2012), “Stochastic Blockmodels with a Growing Number of Classes,” Biometrika, 99, 273–284. 92008Daudin et al.Daudin et al. (2008)daudin2008mixture Daudin, J. J., Picard, F., and Robin, S. (2008), “A Mixture Model for Random Graphs,” Statistics and Computing, 18, 173–183. 102019Di Maggio et al.Di Maggio et al. (2019)di2019relevance Di Maggio, M., Franzoni, F., Kermani, A., and Sommavilla, C. (2019), “The Relevance of Broker Networks for Information Diffusion in the Stock Market,” Journal of Financial Economics, 134, 419–446. 112017Gao et al.Gao et al. (2017)gao2017achieving Gao, C., Ma, Z., Zhang, A. Y., and Zhou, H. H. (2017), “Achieving Optimal Misclassification Proportion in Stochastic Block Models,” The Journal of Machine Learning Research, 18, 1980–2024. 122018Gao et al.Gao et al. (2018)gao2018community Gao, C., Ma, Z., Zhang, A. Y., and Zhou, H. H. (2018), “Community Detection in Degree-Corrected Block Models,” The Annals of Statistics, 46, 2153–2185. 132023Han et al.Han et al. (2023)han2023universal Han, X., Yang, Q., and Fan, Y. (2023), “Universal Rank Inference via Residual Subsampling with Application to Large Networks,” The Annals of Statistics, 51, 1109–1133. 141983Holland et al.Holland et al. (1983)holland1983stochastic Holland, P. W., Laskey, K. B., and Leinhardt, S. (1983), “Stochastic Blockmodels: First Steps,” Social Networks, 5, 109–137. 152021Hu et al.Hu et al. (2021)hu2021using Hu, J., Zhang, J., Qin, H., Yan, T., and Zhu, J. (2021), “Using Maximum Entry-Wise Deviation to Test the Goodness of Fit for Stochastic Block Models,” Journal of the American Statistical Association, 116, 1373–1382. 162016Ji and JinJi and Jin (2016)ji2016coauthorship Ji, P., and Jin, J. (2016), “Coauthorship and Citation Networks for Statisticians,” The Annals of Applied Statistics, 10, 1779–1812. 172015JinJin (2015)jin2015fast Jin, J. (2015), “Fast Community Detection by SCORE,” The Annals of Statistics, 43, 57–89. 182022Jing et al.Jing et al. (2022)jing2022community Jing, B. Y., Li, T., Ying, N., and Yu, X. (2022), “Community Detection in Sparse Networks Using the Symmetrized Laplacian Inverse Matrix (Slim),” Statistica Sinica, 32. 192011Karrer and NewmanKarrer and Newman (2011)karrer2011stochastic Karrer, B., and Newman, M. E. (2011), “Stochastic Blockmodels and Community Structure in Networks,” Physical Review E, 83, 016107. 202016Karwa et al.Karwa et al. (2016)karwa2016exact Karwa, V., Pati, D., Petrovic, S., Solus, L., Alexeev, N., Raic, M., Wilburne, D., Williams, R., and Yan, B. (2016), “Exact Tests for Stochastic Block Models,” arXiv preprint arXiv:1612.06040. 212011Katona et al.Katona et al. (2011)katona2011network Katona, Z., Zubcsek, P. P., and Sarvary, M. (2011), “Network Effects and Personal Influences: The Diffusion of an Online Social Network,” Journal of Marketing Research, 48, 425–443. 222015LawyerLawyer (2015)lawyer2015understanding Lawyer, G. (2015), “Understanding the Influence of All Nodes in a Network,” Scientific Reports, 5, 8665. 232019Lee and WilkinsonLee and Wilkinson (2019)lee2019review Lee, C., and Wilkinson, D. J. (2019), “A Review of Stochastic Block Models and Extensions for Graph Clustering,” Applied Network Science, 4, 1–50. 242016LeiLei (2016)lei2016goodness Lei, J. (2016), “A Goodness-of-Fit Test for Stochastic Block Models,” The Annals of Statistics, 44, 401–424. 252015Lei and RinaldoLei and Rinaldo (2015)lei2015consistency Lei, J., and Rinaldo, A. (2015), “Consistency of Spectral Clustering in Stochastic Block Models,” The Annals of Statistics, 43, 215–237. 262017Lei and ZhuLei and Zhu (2017)lei2017generic Lei, J., and Zhu, L. (2017), “Generic Sample Splitting for Refined Community Recovery in Degree Corrected Stochastic Block Models,” Statistica Sinica, 1639–1659. 272020Li et al.Li et al. (2020)li2020network Li, T., Levina, E., and Zhu, J. (2020), “Network Cross-Validation by Edge Sampling,” Biometrika, 107, 257–276. 282015Mossel et al.Mossel et al. (2015)mossel2015reconstruction Mossel, E., Neeman, J., and Sly, A. (2015), “Reconstruction and Estimation in the Planted Partition Model,” Probability Theory and Related Fields, 162, 431–461. 292018Mossel et al.Mossel et al. (2018)mossel2018proof Mossel, E., Neeman, J., and Sly, A. (2018), “A Proof of the Block Model Threshold Conjecture,” Combinatorica, 38, 665–708. 302018Musumeci et al.Musumeci et al. (2018)musumeci2018overview Musumeci, F., Rottondi, C., Nag, A., Macaluso, I., Zibar, D., Ruffini, M., and Tornatore, M. (2018), “An Overview on Application of Machine Learning Techniques in Optical Networks,” IEEE Communications Surveys &\&& Tutorials, 21, 1383–1408. 312006NewmanNewman (2006)newman2006modularity Newman, M. E. (2006), “Modularity and Community Structure in Networks,” Proceedings of the National Academy of Sciences, 103, 8577–8582. 322001Nowicki and SnijdersNowicki and Snijders (2001)nowicki2001estimation Nowicki, K., and Snijders, T. A. B. (2001), “Estimation and Prediction for Stochastic Blockstructures,” Journal of the American Statistical Association, 96, 1077–1087. 332011Rohe et al.Rohe et al. (2011)rohe2011spectral Rohe, K., Chatterjee, S., and Yu, B. (2011), “Spectral Clustering and the High-Dimensional Stochastic Blockmodel,” The Annals of Statistics, 39, 1878–1915. 342017Saldana et al.Saldana et al. (2017)saldana2017many Saldana, D. F., Yu, Y., and Feng, Y. (2017), “How Many Communities Are There?,” Journal of Computational and Graphical Statistics, 26, 171–181. 352015Sarkar and BickelSarkar and Bickel (2015)sarkar2015role Sarkar, P., and Bickel, P. J. (2015), “Role of Normalization in Spectral Clustering for Stochastic Blockmodels,” The Annals of Statistics, 43, 962–990. 361997Snijders and NowickiSnijders and Nowicki (1997)snijders1997estimation Snijders, T. A., and Nowicki, K. (1997), “Estimation and Prediction for Stochastic Blockmodels for Graphs with Latent Block Structure,” Journal of Classification, 14, 75–100. 372018Tabassum et al.Tabassum et al. (2018)tabassum2018social Tabassum, S., Pereira, F. S., Fernandes, S., and Gama, J. (2018), “Social Network Analysis: An Overview,” Wiley Interdisciplinary Reviews: Data Mining and Knowledge Discovery, 8, e1256. 382010ValenteValente (2010)valente2010social Valente, T. W. (2010), Social Networks and Health: Models, Methods, and Applications, Oxford University Press. 392017Wang and BickelWang and Bickel (2017)wang2017likelihood Wang, Y. R., and Bickel, P. J. (2017), “Likelihood-Based Model Selection for Stochastic Block Models,” The Annals of Statistics, 45, 500–528. 402011Westveld and HoffWestveld and Hoff (2011)westveld2011mixed Westveld, A. H., and Hoff, P. D. (2011), “A Mixed Effects Model for Longitudinal Relational and Network Data, with Applications to International Trade and Conflict,” The Annals of Applied Statistics, 5, 843–872. 412022Wu et al.Wu et al. (2022)wu2022inward Wu, Y., Lan, W., Zou, T., and Tsai, C. L. (2022), “Inward and Outward Network Influence Analysis,” Journal of Business &\&& Economic Statistic, 40, 1617–1628. 422016Zhang and ZhouZhang and Zhou (2016)zhang2016minimax Zhang, A. Y., and Zhou, H. H. (2016), “Minimax Rates of Community Detection in Stochastic Block Models,” The Annals of Statistics, 44, 2522–2280. 432023Zhang and AminiZhang and Amini (2023)zhang2020adjusted Zhang, L., and Amini, A. A. (2023), “Adjusted Chi-Square Test for Degree-Corrected Block Models,” The Annals of Statistics, 51, 2366–2385. 442012Zhao et al.Zhao et al. (2012)zhao2012consistency Zhao, Y., Levina, E., and Zhu, J. (2012), “Consistency of Community Detection in Networks Under Degree-Corrected Stochastic Block Models,” The Annals of Statistics, 40, 2266–2292.

### v2

@@ 4 Real data analyses
To demonstrate the usefulness of our proposed method for both dense and sparse networks, we analyze two real datasets. Based on simulation studies, we find that our proposed method and Hu et al. (2021)’s method generally outperform those methods proposed by Lei (2016), Han et al. (2023), and Zhang and Amini (2023). Thus, in our real data analyses, we only consider our proposed method and Hu et al. (2021)’s method. The first dataset is the international trade data, which has been studied by Westveld and Hoff (2011), Lei (2016), and Hu et al. (2021). This data set contains yearly trade information of 58 countries. Hu et al. (2021) constructed a network with density of 48.1% (=1590/(58×57)×1590/(58$\times$ 57$\times$1590 / ( 58 × 57 ) × 100%), and it is a typical dense network. As noticed by Hu et al. (2021), the augmentation procedure may not work well for this data set because the size of the network is small and the added community may have less than or equal to two nodes. Therefore, we apply the test statistics, $\Theta_{n,boot}$ and $\Theta_{Hu,boot}$, to analyze the international trade network. The second data set is the co-authorship network dataset collected by Ji and Jin (2016). This network has 3,607 nodes and the density of this network is approximately equal to 0.012% (=1561/(3607×3606)×1561/(3607$\times$ 3606$\times$1561 / ( 3607 × 3606 ) ×100%), which makes it a sparse network. The simulation results indicate that $\Theta^{+}_{n,boot}$ and $\Theta^{+}_{Hu,boot}$ perform better than $\Theta_{n}$ and $\Theta_{Hu}$, respectively. Hence, we employ them to study the co-authorship network. However, the augmented tests are not applicable for testing $H_{0}:K_{0}=1$ since it is infeasible to specify the connection probability between communities. Thus, $\Theta_{n,boot}$ and $\Theta_{Hu,boot}$ are considered in this case. Finally, the number of bootstraps used in both examples is $M=100$, and the number of samples drawn from $n$ nodes is $B=\tilde{Q}^{-1/2}\{n/\log n\}^{1/3}$ as given in simulation studies.
@@ 4.1 Dense network case: International trade data
This dataset contains yearly trade information for $n=58$ countries from 1981 to 2000. For any given year, $\text{Trade}_{ij}$ represents the value of exports from country $i$ to country $j$; further descriptions can be found in Westveld and Hoff (2011). Following Saldana et al. (2017) and Hu et al. (2021), we consider the international trade data in 1995, and construct the network from the 58 countries based on $\text{Trade}_{ij}$ as follows. Each node represents a country, and define $W_{ij}=\text{Trade}_{ij}+\text{Trade}_{ji}$. Then, the adjacency matrix $A=(a_{ij})$ is defined such that $a_{ij}=1$ if $W_{ij}>W_{0.5}$ and $a_{ij}=0$ otherwise, where $W_{0.5}$ is the 50% quantile of the $W_{ij}$ values. As discussed in Saldana et al. (2017), it is possible to decompose the network into 3, 7, and 10 communities by three different methods. However, for different choice of $K$, the resulting communities from these 58 countries vary. Therefore, determining the appropriate number of communities is critical to label communities and make meaningful interpretations for 58 countries. To this end, we employ the statistics $\Theta_{n,boot}$ and $\Theta_{Hu,boot}$ to test the three hypotheses $H_{0}:K=3$, $H_{0}:K=7$, and $H_{0}:K=10$, respectively.
Given the 5% significance level, we apply our Theorem 1 and Hu et al. (2021)’s Theorem 1 to obtain the critical values of $\Theta_{n,boot}$ and $\Theta_{Hu,boot}$, which are $t_{n,0.95}=4.792$ and $t_{Hu,0.95}=3.413$, respectively. For $K_{0}=3,7,$ and 10, Hu et al. (2021) calculated their corresponding test statistics, $\Theta_{Hu,boot}=44.28,3.33$ and 13.44. Comparing them with the critical value, 3.413, they concluded that $K=7$ is a reasonable choice. We next compute our proposed test statistics, $\Theta_{n,boot}=15.02,3.07,$ and 16.76, for $K_{0}=3,7,$ and 10, respectively. Comparing them with the critical value, 4.792, we find that $\Theta_{n,boot}$ does not reject the hypothesis $H_{0}:K=7$, whereas it rejects the hypotheses $H_{0}:K=3$ and $H_{0}:K=10$. Hence, $\Theta_{n,boot}$ yields the same finding as that of Hu et al. (2021).
@@ 4.2 Sparse network case: Co-authorship data
Using this sparse network data, we test the hypothetical number of communities $K_{0}\in[1,8]$. Given the 5% significance level, we have $t_{n,0.95}=4.792$ and $t_{Hu,0.95}=3.413$ as given in the above example. In testing $H_{0}:K=1$, both $\Theta_{n,boot}$ and $\Theta_{Hu,boot}$ are larger than 400, which results in rejecting the null hypothesis. For $K_{0}\in[2,8]$, we have $\Theta^{+}_{Hu,boot}>7$, which results in rejecting all of their corresponding hypotheses. This finding is not surprising since $\Theta^{+}_{Hu,boot}$ fails to converge to an extreme-value distribution for sparse networks. In testing $K_{0}\in[2,8]$, we calculate $\Theta^{+}_{n,boot}$=14.81, 15.80, 4.23, -1.53, 3.75, -2.52, and 3.08. As a result, $\Theta^{+}_{n,boot}$ does not reject each of those hypotheses from $H_{0}:K=4$ to $H_{0}:K=8$ at the 5% significance level. Hence, $\Theta^{+}_{n,boot}$ can identify $K=4$ to $K=8$ communities for all authors.
@@
In sum, our proposed test and Hu et al. (2021)’s test lead to the same result in the dense network example, and we also provide more interpretations. In the sparse network example, our test successfully detects a reasonable number of network communities, whereas Hu et al. (2021)’s test rejects all possible hypotheses and fails to yield any conclusion. Consequently, our proposed method is applicable for both data sets.
@@ 5 Concluding remarks
12013Agarwal et al.Agarwal et al. (2013)agarwal2013traffic Agarwal, S., Kodialam, M., and Lakshman, T. (2013), “Traffic Engineering in Software Defined Networks,” 2013 Proceedings IEEE INFOCOM, 2211–2219. 22013Amini et al.Amini et al. (2013)amini2013pseudo Amini, A. A., Chen, A., Bickel, P. J., and Levina, E. (2013), “Pseudo-Likelihood Methods for Community Detection in Large Sparse Networks,” The Annals of Statistics, 41, 2097–2122. 32016Andrikopoulos et al.Andrikopoulos et al. (2016)andrikopoulos2016four Andrikopoulos, A., Samitas, A., and Kostaris, K. (2016), “Four Decades of the Journal of Econometrics: Coauthorship Patterns and Networks,” Journal of Econometrics, 195(1), 23–32. 42009Bickel and ChenBickel and Chen (2009)bickel2009nonparametric Bickel, P. J., and Chen, A. (2009), “A Nonparametric View of Network Models and Newman–Girvan and Other Modularities,” Proceedings of the National Academy of Sciences, 106, 21068–21073. 52016Bickel and SarkarBickel and Sarkar (2016)bickel2016hypothesis Bickel, P. J., and Sarkar, P. (2016), “Hypothesis Testing for Automated Community Detection in Networks,” Journal of the Royal Statistical Society Series B: Statistical Methodology, 78, 253–273. 62009Bramoullé et al.Bramoullé et al. (2009)bramoulle2009identification Bramoullé, Y., Djebbari, H., and Fortin, B. (2009), “Identification of Peer Effects through Social Networks,” Journal of Econometrics, 150(1), 41–55. 72015ChatterjeeChatterjee (2015)chatterjee2015matrix Chatterjee, S. (2015), “Matrix Estimation by Universal Singular Value Thresholding,” The Annals of Statistics, 43, 177–214. 82018Chen and LeiChen and Lei (2018)chen2018network Chen, K., and Lei, J. (2018), “Network Cross-Validation for Determining the Number of Communities in Network Data,” Journal of the American Statistical Association, 113, 241–251. 92015Chin et al.Chin et al. (2015)chin2015stochastic Chin, P., Rao, A., and Vu, V. (2015), “Stochastic Block Model and Community Detection in Sparse Graphs: A Spectral Algorithm with Optimal Rate of Recovery,” Conference on Learning Theory, 391–423. 102012Choi et al.Choi et al. (2012)choi2012stochastic Choi, D. S., Wolfe, P. J., and Airoldi, E. M. (2012), “Stochastic Blockmodels with a Growing Number of Classes,” Biometrika, 99, 273–284. 112008Daudin et al.Daudin et al. (2008)daudin2008mixture Daudin, J. J., Picard, F., and Robin, S. (2008), “A Mixture Model for Random Graphs,” Statistics and Computing, 18, 173–183. 122019Di Maggio et al.Di Maggio et al. (2019)di2019relevance Di Maggio, M., Franzoni, F., Kermani, A., and Sommavilla, C. (2019), “The Relevance of Broker Networks for Information Diffusion in the Stock Market,” Journal of Financial Economics, 134, 419–446. 132025Fu et al.Fu et al. (2025)Fu2025twosample Fu, K., Hu, J., Keita, S., and Liu, H. (2025), “Two-Sample Test for Stochastic Block Models Via Maximum Entry-Wise Deviation,” Journal of the Korean Statistical Society, 18, 299–313. 142017Gao et al.Gao et al. (2017)gao2017achieving Gao, C., Ma, Z., Zhang, A. Y., and Zhou, H. H. (2017), “Achieving Optimal Misclassification Proportion in Stochastic Block Models,” The Journal of Machine Learning Research, 18, 1980–2024. 152018Gao et al.Gao et al. (2018)gao2018community Gao, C., Ma, Z., Zhang, A. Y., and Zhou, H. H. (2018), “Community Detection in Degree-Corrected Block Models,” The Annals of Statistics, 46, 2153–2185. 162016Härdle et al.Härdle et al. (2016)hardle2016tenet Härdle, W. K., Wang, W., and Yu, L. (2016), “TENET: Tail-Event Driven Network Risk,” Journal of Econometrics, 192(2), 499–513. 172023Han et al.Han et al. (2023)han2023universal Han, X., Yang, Q., and Fan, Y. (2023), “Universal Rank Inference via Residual Subsampling with Application to Large Networks,” The Annals of Statistics, 51, 1109–1133. 181983Holland et al.Holland et al. (1983)holland1983stochastic Holland, P. W., Laskey, K. B., and Leinhardt, S. (1983), “Stochastic Blockmodels: First Steps,” Social Networks, 5, 109–137. 192021Hu et al.Hu et al. (2021)hu2021using Hu, J., Zhang, J., Qin, H., Yan, T., and Zhu, J. (2021), “Using Maximum Entry-Wise Deviation to Test the Goodness of Fit for Stochastic Block Models,” Journal of the American Statistical Association, 116, 1373–1382. 202016Ji and JinJi and Jin (2016)ji2016coauthorship Ji, P., and Jin, J. (2016), “Coauthorship and Citation Networks for Statisticians,” The Annals of Applied Statistics, 10, 1779–1812. 212015JinJin (2015)jin2015fast Jin, J. (2015), “Fast Community Detection by SCORE,” The Annals of Statistics, 43, 57–89. 222024Jin et al.Jin et al. (2024)jin2024mixed Jin, J., Ke, Z. T., and Luo, S. (2024), “Mixed Membership Estimation for Social Networks,” Journal of Econometrics, 239(2), 105369. 232022Jing et al.Jing et al. (2022)jing2022community Jing, B. Y., Li, T., Ying, N., and Yu, X. (2022), “Community Detection in Sparse Networks Using the Symmetrized Laplacian Inverse Matrix (Slim),” Statistica Sinica, 32. 242024JochmansJochmans (2024)jochmans2024nonparametric Jochmans, K. (2024), “Nonparametric Identification and Estimation of Stochastic Block Models from Many Small Networks,” Journal of Econometrics, 242(2), 105805. 252011Karrer and NewmanKarrer and Newman (2011)karrer2011stochastic Karrer, B., and Newman, M. E. (2011), “Stochastic Blockmodels and Community Structure in Networks,” Physical Review E, 83, 016107. 262016Karwa et al.Karwa et al. (2016)karwa2016exact Karwa, V., Pati, D., Petrovic, S., Solus, L., Alexeev, N., Raic, M., Wilburne, D., Williams, R., and Yan, B. (2016), “Exact Tests for Stochastic Block Models,” arXiv preprint arXiv:1612.06040. 272011Katona et al.Katona et al. (2011)katona2011network Katona, Z., Zubcsek, P. P., and Sarvary, M. (2011), “Network Effects and Personal Influences: The Diffusion of an Online Social Network,” Journal of Marketing Research, 48, 425–443. 282019Lee and WilkinsonLee and Wilkinson (2019)lee2019review Lee, C., and Wilkinson, D. J. (2019), “A Review of Stochastic Block Models and Extensions for Graph Clustering,” Applied Network Science, 4, 1–50. 292016LeiLei (2016)lei2016goodness Lei, J. (2016), “A Goodness-of-Fit Test for Stochastic Block Models,” The Annals of Statistics, 44, 401–424. 302015Lei and RinaldoLei and Rinaldo (2015)lei2015consistency Lei, J., and Rinaldo, A. (2015), “Consistency of Spectral Clustering in Stochastic Block Models,” The Annals of Statistics, 43, 215–237. 312017Lei and ZhuLei and Zhu (2017)lei2017generic Lei, J., and Zhu, L. (2017), “Generic Sample Splitting for Refined Community Recovery in Degree Corrected Stochastic Block Models,” Statistica Sinica, 1639–1659. 322020Li et al.Li et al. (2020)li2020network Li, T., Levina, E., and Zhu, J. (2020), “Network Cross-Validation by Edge Sampling,” Biometrika, 107, 257–276. 332015Mossel et al.Mossel et al. (2015)mossel2015reconstruction Mossel, E., Neeman, J., and Sly, A. (2015), “Reconstruction and Estimation in the Planted Partition Model,” Probability Theory and Related Fields, 162, 431–461. 342018Mossel et al.Mossel et al. (2018)mossel2018proof Mossel, E., Neeman, J., and Sly, A. (2018), “A Proof of the Block Model Threshold Conjecture,” Combinatorica, 38, 665–708. 352018Musumeci et al.Musumeci et al. (2018)musumeci2018overview Musumeci, F., Rottondi, C., Nag, A., Macaluso, I., Zibar, D., Ruffini, M., and Tornatore, M. (2018), “An Overview on Application of Machine Learning Techniques in Optical Networks,” IEEE Communications Surveys &\&& Tutorials, 21, 1383–1408. 362006NewmanNewman (2006)newman2006modularity Newman, M. E. (2006), “Modularity and Community Structure in Networks,” Proceedings of the National Academy of Sciences, 103, 8577–8582. 372001Nowicki and SnijdersNowicki and Snijders (2001)nowicki2001estimation Nowicki, K., and Snijders, T. A. B. (2001), “Estimation and Prediction for Stochastic Blockstructures,” Journal of the American Statistical Association, 96, 1077–1087. 382011Rohe et al.Rohe et al. (2011)rohe2011spectral Rohe, K., Chatterjee, S., and Yu, B. (2011), “Spectral Clustering and the High-Dimensional Stochastic Blockmodel,” The Annals of Statistics, 39, 1878–1915. 392017Saldana et al.Saldana et al. (2017)saldana2017many Saldana, D. F., Yu, Y., and Feng, Y. (2017), “How Many Communities Are There?,” Journal of Computational and Graphical Statistics, 26, 171–181. 402015Sarkar and BickelSarkar and Bickel (2015)sarkar2015role Sarkar, P., and Bickel, P. J. (2015), “Role of Normalization in Spectral Clustering for Stochastic Blockmodels,” The Annals of Statistics, 43, 962–990. 411997Snijders and NowickiSnijders and Nowicki (1997)snijders1997estimation Snijders, T. A., and Nowicki, K. (1997), “Estimation and Prediction for Stochastic Blockmodels for Graphs with Latent Block Structure,” Journal of Classification, 14, 75–100. 422017Wang and BickelWang and Bickel (2017)wang2017likelihood Wang, Y. R., and Bickel, P. J. (2017), “Likelihood-Based Model Selection for Stochastic Block Models,” The Annals of Statistics, 45, 500–528. 432011Westveld and HoffWestveld and Hoff (2011)westveld2011mixed Westveld, A. H., and Hoff, P. D. (2011), “A Mixed Effects Model for Longitudinal Relational and Network Data, with Applications to International Trade and Conflict,” The Annals of Applied Statistics, 5, 843–872. 442022Wu et al.Wu et al. (2022)wu2022inward Wu, Y., Lan, W., Zou, T., and Tsai, C. L. (2022), “Inward and Outward Network Influence Analysis,” Journal of Business &\&& Economic Statistic, 40, 1617–1628. 452024aWu et al.Wu et al. (2024a)wu2024atwosample Wu, Q., and Hu, J. (2024a), “Two-Sample Test of Stochastic Block Models,” Computational Statistics &\&& Data Analysis, 192, 107903. 462024Wu et al.Wu et al. (2024)wu2024twosample Wu, Q., and Hu, J. (2024), “Two-Sample Test of Stochastic Block Models Via the Maximum Sampling Entry-Wise Deviation,” Journal of the Korean Statistical Society, 53, 617–636. 472016Zhang and ZhouZhang and Zhou (2016)zhang2016minimax Zhang, A. Y., and Zhou, H. H. (2016), “Minimax Rates of Community Detection in Stochastic Block Models,” The Annals of Statistics, 44, 2522–2280. 482023Zhang and AminiZhang and Amini (2023)zhang2020adjusted Zhang, L., and Amini, A. A. (2023), “Adjusted Chi-Square Test for Degree-Corrected Block Models,” The Annals of Statistics, 51, 2366–2385. 492012Zhao et al.Zhao et al. (2012)zhao2012consistency Zhao, Y., Levina, E., and Zhu, J. (2012), “Consistency of Community Detection in Networks Under Degree-Corrected Stochastic Block Models,” The Annals of Statistics, 40, 2266–2292.